# MoMa: Efficient Early-Fusion Pre-training with Mixture of Modality-Aware Experts

### v1

@@ 1 Introduction
The insight driving our approach is the inherent heterogeneity of modalities: text and image tokens possess distinct information densities and redundancy patterns (Liang et al., 2023). While we integrate these tokens into a unified early-fusion architecture, we propose further optimizing this framework by incorporating modality-specific modules. This concept, which we term modality-aware sparsity (MaS), enables models to better capture features specific to each modality while maintaining strong cross-modality integration through partial parameter sharing and attention mechanisms. Previous work such as VLMo (Bao et al., 2022), BEiT-3 (Wang et al., 2022a) and VL-MoE (Shen et al., 2023) have applied mixture-of-modality-experts (MoME) to train vision-language encoders and masked language models. We apply this type of sparse component to the mixed-modal, early-fusion language modeling setting.
@@ 5 Related Work
Recent advancements in sparse modeling techniques have shown promising results in scaling language models efficiently. VL-MoE (Shen et al., 2023) proposed a vision-language model that uses a sparse mixture-of-experts (MoE) to scale efficiently. It applies sparse MoE to the vision and language feed-forward networks in a unified framework, using masked data modeling objectives for pretraining. VL-MoE achieves strong performance on vision-language tasks while using fewer parameters per token than dense models, providing additional evidence for MoE training for multi-modal settings.

### v2

@@ 1 Introduction
The insight driving our approach is the inherent heterogeneity of modalities: text and image tokens possess distinct information densities and redundancy patterns (Liang et al., 2023). While we integrate these tokens into a unified early-fusion architecture, we propose further optimizing this framework by incorporating modality-specific modules. This concept, which we term modality-aware sparsity (MaS), enables models to better capture features specific to each modality while maintaining strong cross-modality integration through partial parameter sharing and attention mechanisms. Previous work such as VLMo (Bao et al., 2022), BEiT-3 (Wang et al., 2022a) and VL-MoE (Shen et al., 2023) have applied mixture-of-modality-experts to train vision-language encoders and masked language models. We apply this type of sparse component to the mixed-modal, early-fusion language modeling setting.
@@ 5 Related Work
Sparse neural networks.
Sparse neural networks have emerged as a promising approach to improve the efficiency and scalability of deep learning models. One notable architecture is the Mixture-of-Experts (MoE) model, which dynamically selects a subset of experts to process each input, reducing computational costs and promoting specialization (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022). These advances in sparse neural networks have paved the way for the development of increasingly large and powerful models, enabling state-of-the-art results in various domains, including natural language processing (Jiang et al., 2024) and computer vision (Riquelme et al., 2021). Recent work have also proposed approaches that trains mixture-of-experts network using dense model as the seed to improve training stability (Komatsuzaki et al., 2023) and leverage efficiency gains from asynchronous computation (Sukhbaatar et al., 2024).
@@
Recent advancements in sparse modeling techniques have also shown promising results in efficient scaling of multimodal language models. VL-MoE (Shen et al., 2023) proposed a vision-language model that uses a sparse mixture-of-experts (MoE) to scale efficiently. It applies modality-aware MoE (Bao et al., 2022; Wang et al., 2022a; Shen et al., 2023; Chen et al., 2024) to the vision and language feed-forward networks in a unified framework, using masked data modeling objectives for pretraining. VL-MoE achieves strong performance on vision-language tasks while using fewer parameters per token than dense models, providing additional evidence for MoE training for multi-modal settings.